In 431 NFL football games that went to overtime, the teams that won the coin toss...

  1. In 431 NFL football games that went to overtime, the teams that won the coin toss went on to win 235 of those games. If the coin-toss method is fair, we expect that the teams winning the coin would win about 50% of the games, so we expect about 215.5 wins in 431 overtime games. Assuming that there is a0.5 probability of winning a game after winning the coin toss, find the probability of getting at least 235 winning games among the 431 teams that won the coin toss. That is, given n = 431 and p = 0.5, find the probability of at least 235 wins.
  1. 0.3803
  2. 0.0035
  3. 0.9664
  4. 0.0336                                                                                  E. None
  1. Identify the population: A survey of 500 adults in the U.S. found that 54% drink coffee daily.
  1. Collection of the 500 adults surveyed
  2. Collection of all adults in the U.S.
  3. 54%
  4. 500                                                                                        E. None
  1. Identify the sample: A survey of 500 adults in the U.S. found that 54% drink coffee daily.
  1. Collection of the 500 adults surveyed
  2. Collection of all adults in the U.S.
  3. 54%
  4. 500                                                                                       E. None
  1. Identify the data set’s level of measurement: The IQ scores of students in a class.
  1. Nominal
  2. Ordinal
  3. Interval
  4. Ratio                                                                                    E. None
  1. Identify the data set’s level of measurement: The nationality of each person on an airplane.
  1. Nominal
  2. Ordinal
  3. Interval
  4. Ratio                                                                                   E. None

  1. Identify the data set’s level of measurement: The salaries of nurses at a hospital.
  1. Nominal
  2. Ordinal
  3. Interval
  4. Ratio                                                                                    E. None
  1. Decide which method of data collection would be most appropriate: A study of the effect of using MyStatLab (as an application tool) on grades in a statistics course.
  1. Observational study
  2. Experiment
  3. Simulation
  4. Survey                                                                                  E. None
Answer #1

In 431 NFL football games that went to overtime, the teams that won the coin toss went on to win 235 of those games. If the coin-toss method is fair, we expect that the teams winning the coin would win about 50% of the games, so we expect about 215.5 wins in 431 overtime games. Assuming that there is a0.5 probability of winning a game after winning the coin toss, find the probability of getting at least 235 winning games among the 431 teams that won the coin toss. That is, given n = 431 and p = 0.5, find the probability of at least 235 wins.

correct answer D. 0.0336

Mean = np= 431*0.5=215.5

standard deviation = sqrt(npq)=sqrt(431*0.5*0.5)=10.38027

We want the probability of at least 235 wins, so x = 235.

Z value=(x-mean)/standard deviation = (235-215.5)/10.38027 = 1.83

The probability is 0.0336 (Ref Z-Table) for the coin flip winning team to win at least 235 games.

Identify the population: A survey of 500 adults in the U.S. found that 54% drink coffee daily.

correct answer B. Collection of all adults in the U.S.

Identify the sample: A survey of 500 adults in the U.S. found that 54% drink coffee daily.

correct answer A.Collection of the 500 adults surveyed

Identify the data set’s level of measurement: The IQ scores of students in a class.

correct answer C. Interval